OpenStudy (anonymous):

is (2,0) the answer to this system below y=2x-4 x+3y=5

OpenStudy (anonymous):

use substitution to substitute the y in for x+3(2x-4)=5

OpenStudy (anonymous):

3x+12=5?

OpenStudy (anonymous):

no you have to distribute the 3 to 2x and -4 to get x+6x-4=5

OpenStudy (texaschic101):

check your math Michiganboy

OpenStudy (anonymous):

wait -12 sorry

OpenStudy (anonymous):

x+6x-12=5

OpenStudy (anonymous):

then just solve for x

OpenStudy (anonymous):

so its not a solution i got 17/7

OpenStudy (anonymous):

x+6x-12=5 7x-12=5 +12 +12 ------------ 7x=17 --- --- 7 7 x= 17/7 your right :P

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem